A PRODUCT DECOMPOSITION OF Ω_{0}^{3}ΣRP^{2} Jie Wu

Abstract. We give a specific product decomposition of the base-point path connected component of the triple loop space of the suspension of the projective plane.

Typeset by AMS-TEX